# A Survey on Causal Discovery: Theory and Practice

### v1

@@ 2.7 Adding Prior Knowledge
[…]
@@ 3.1 Constraint-based Algorithms
The null $H_{0}$ and alternative hypotheses $H_{1}$ defined as $H_{0}:X\perp\!\!\!\perp_{P}Y\,|\,\textbf{Z}$ and $H_{1}:X\not\perp\!\!\!\perp_{P}Y\,|\,\textbf{Z}$, let $I(X,Y|\textbf{Z})$ to denote a conditional independence (CI) test. The null hypothesis $H_{0}$ is not rejected if and only if the resulting p-value is higher than a chosen significance level $\alpha$:
@@
When faithfulness is assumed, probability independence implies graphical separation [^2] Here the term separation is used as a placeholder for a generic graphical separation, which is intended as d-separation for directed graphs and m-separation for mixed graphs.. The main limitation of this approach is related to the exponential growth of the conditioning set Z. Indeed, given the pair $(X,Y)$, in the worst case scenario where $X$ is dependent on $Y$ (or vice-versa), the algorithm is required to test for $\big{|}2^{(\textbf{V}\backslash\{X,Y\})}\big{|}$ conditioning sets.
@@
Conditional Independence with Mixed Data
Constrain-based techniques are essentially agnostic of the specific conditional independence test that is being used. Indeed, it is possible to take advantage of such approaches in a wide variety of scenarios, as long as the assumptions of the said test are satisfied. While the main focus of causal discovery studies has been into either discrete or continuous settings, recent advances in conditional independence testing [6, 107] extend existing tests to mixed-data.
Peter-Clark (PC)
One of the most studied algorithm that leverages the CISs is the Peter-Clark (PC) algorithm [99] with its variants [18, 49].
The first step of the procedure consists in defining a complete undirected graph over the variables of the given data set D. Subsequently, a sequence of conditional independence (CI) tests are performed following an heuristic strategy [18], in order to minimize the number of tests needed. For instance, it is known that the power of CI test decreases when the size of the conditioning set increases [51], due to the curse of dimensionality. A common approach consists in selecting an upper limit to the size of the conditioning set, discarding computational-intensive time-wasting tests with low significance levels.
The obtained independence statements are then used to remove the associated edges and identify the underlying skeleton. Finally, the remaining edges are oriented according to a set of rules [58] that leverages the identified v-structures and acyclicity property.
@@
Fast Causal Inference (FCI)
A first extension of the PC algorithm to the causal insufficient setting (Subsection 2.6) is represented by the Fast Causal Inference (FCI) [102, 113] algorithm. Specifically, the FCI algorithm relaxes both the assumption of no latent confounding [32] and no selection bias [50] in the observational setting, pushing the causal discovery problem a step closer to real-world scenarios. In this context, the authors leverage the definition of discriminating path to derive a new set of orientation rules.
@@
Let $G$ be an ancestral graph, a path $\pi=(X,\dots,W,Z,Y)$ between $X$ and $Y$ is a discriminating path for $Z$ if (i) $\pi$ contains at least three edges, (ii) $X$ is not adjacent to $Y$, (iii) $Z$ is adjacent to $Y$, and (iv) every vertex between $X$ and $Z$ is a collider on $\pi$ and parent of $Y$.
@@ 3.2 Score-based Algorithms
In the next few paragraphs, a set of properties for scoring criteria are introduced, before shifting the focus on an optimal two-step procedure for the causal sufficient scenario.
@@
Let D be a data set associated with a probability distribution $P$ and let $G$ and $H$ be two graphs. A scoring criterion $\mathcal{S}$ is said to be consistent in the limit of the number of samples if and only if:
@@
  • $X\perp\!\!\!\perp_{P}Y\,|\,Pa(X)\implies\mathcal{S}(G,\textbf{D})<\mathcal{S}(H,\textbf{D})$,
  • $X\not\perp\!\!\!\perp_{P}Y\,|\,Pa(X)\implies\mathcal{S}(G,\textbf{D})>\mathcal{S}(H,\textbf{D})$.
@@
This property guarantees that any deletion of an unnecessary edge will produce a higher score value, allowing the definition of a optimal greedy search algorithm.
We report a brief list of scores for reference, such as the Akaike Information Criterion (AIC) [3], the Bayesian Information Criterion (BIC) [87], the Bayesian Dirichlet equivalent uniform (BDeu) [26] and the Bayesian Dirichlet sparse (BDs) [89].
@@
Let $[G]^{*}$ be the equivalence class that is a perfect map of the probability distribution $P$ and D the associated data set. In the limit of the number of samples:
@@
Greedy Equivalent Search (GES)
The Greedy Equivalence Search (GES) [4, 57] is optimal in the limit of the number of samples [17]. The first step of the algorithm consists in the initialization of the empty graph $G$. The algorithm is composed by two phases: the forward search and the backward search. In the forward search phase, (i) $G$ is modified by repeatedly adding the edge that has the highest delta score, until there is no such edge that increases the score. In the backward search phase, (ii) the edge that again achieves the highest delta score is repeatedly removed. The algorithm terminates once it reaches a local maximum during the backward search phase.
@@
Fast GES (FGES)
Score-based algorithms are as fast as the computation of the chosen scoring criterion is. Leveraging the properties of the score function, it is possible to minimize the number of computations needed by storing previous intermediate evaluations. Not only this optimizations reduce the computation time considerably, but also allow the application of these methods to high-dimensional data sets [6, 75]. This “fast” variant of GES (FGES) caches partial graph scores (i.e. delta scores), significantly increasing the memory usage, since relevant fragments of the graph may be considered. Moreover, computationally expensive sections of the algorithm can be parallelized, taking advantage of high performance computing (HPC) settings.
@@ 3.3 Hybrid Algorithms
Adaptively Restricted GES (ARGES)
@@
The novelty of this hybrid version of GES stems from the concept of admissible edge. Let $G$ be a CPDAG and $X$ and $Y$ be a pair of non adjacent vertices on it. Adding an edge between $X$ and $Y$ is admissible for the graph $G$ if (i) $X$ and $Y$ are adjacent in the (estimated) skeleton of $G$ or (ii) there exists a node $Z$ such that $X\rightarrow Z\leftarrow Y$ is a v-structure in $G$.
@@
Greedy FCI (GFCI)
Score-based causal discovery algorithms such as GES and FGES are asymptotically correct, but are not designed to work in a causal insufficient scenario, where unmeasured confounders are present in the true graph. Constraint-based causal search algorithms, such as FCI, are asymptotically correct even with unmeasured confounders, but often perform poorly on small samples. The Greedy Fast Causal Inference (GFCI) [69] algorithm combines score-based and constraint-based algorithms improving over the previous results while being asymptotically correct under causal insufficiency.
@@ 3.4 Other Methods
Hidden Compact Representation (HCR)
@@
The authors address the problem of identifiability of the model and prove that, under mild conditions, the causal graph recovered from observational data is identifiable. The method is tested against both synthetic and real-world data, providing reference values for performance metrics. In these experiments, HCR outperforms linear models in terms of accuracy and sensitivity, especially when the additive noise assumption does not hold.
Quantile Causal Discovery (bQCD)
The quantile causal discovery (bQCD) [105] technique is designed to uncover cause-effect pairs in the bivariate setting. By re-expressing independence statements in light of the minimum description length (MDL) [81], the authors build a discovery procedure by using quantile scoring.
Following [41], let $X$ and $Y$ be two random variables with joint, marginal and conditional distributions denoted by $F$, $F_{X}$ and $F_{X|Y}$ respectively. The key concept here is that a lower complexity follows from a correct causal orientation of the $(X,Y)$ pair, since it is a more informative representation of the associated data.
Hence, the Kolmogorov complexity $K(F)$ is defined as the length of the shortest program that outputs $F(X)$. Since $K(F)$ measures the information contained in $F$, authors in [104] state that if $X$ causes $Y$, then $K(F_{X})+K(F_{Y|X})\leq K(F_{Y})+K(F_{X|Y})$. The problem is that $K(F)$ cannot be computed in practice. Therefore, the authors relies on the MDL principle as a proxy for the Kolmogorov complexity. Such an approximation can be performed by estimating the population quantiles through nonparametric quantile regression.
The resulting procedure is robust to outliers and can be generalized to a wide range of distributions, although it requires that all population quantiles must be computable, which could be a limiting factor in real-world applications.
Linear Non-Gaussian Acyclic (LiNGAM)
@@
The LiNGAM [94, 95] algorithm relies on independent component analysis (ICA) [20] to identify a possible solution for A. In fact, multiple mixing matrices A are feasible solutions for the given joint probability distribution. This technique is essentially focused on discovering asymmetries in the sample distribution to determine the correct causal ordering. Once such ordering has been discovered, the causal graph is built by recovering all and only the edges coherent with the order.
The LiNGAM method has been extended later for causally insufficient settings [34]. Let f be the vector of latent variables and $\boldsymbol{\Lambda}$ the matrix of the connections strength between f and x, then:
@@
The LiNGAM algorithm consistently estimates the connection matrix B. While standard ICA does not scale well in high-dimensional settings, approximated variants of ICA can be used to compute the components with a predefined fix number of iterations with reasonable precision. This leads to a efficient solution in presence of non-gaussian noise and causally insufficient data sets.
Continuous Optimization (NOTEARS)
In the “DAGs with NO TEARS” [114] algorithm, the causal discovery problem is reduced to a continuous optimization problem. The acyclicity constraint is expressed as an equality constraint $h(\textbf{W})=0$, where $h$ is a smooth differentiable function that measures the “DAG-ness” (i.e. a quantification of the acyclicity violations) of a given adjacency matrix W. When W is binary then:
@@
Coefficients smaller than a fixed threshold $\omega>0$ are set to zero, rounding the solution with an arbitrary precision. The evaluation of the matrix exponential is $O(n^{3})$, i.e. cubic in the number of vertices. Given the low computational complexity, NOTEARS outperforms existing methods when both the in-degree and the sample size are large.
@@ 4.1 Cyclic SCM
In a SCM, the causal graph induces a functional set F where equations follows the decomposition enforced by the causal edge assumption, Subsection 2.9. If the causal graph is acyclic, then the SCM itself is called acyclic, or recursive SEM. The concept of recursion is linked to the hierarchical order that arises from the topological ordering of the underlying DAG. Indeed, it is possible to define a sequence $X_{1},X_{2},\dots,X_{n}$ of vertices over V such that for any $X_{i}$ and $X_{j}$ where $i<j$, $X_{j}$ is not a cause of $X_{i}$ [9].
@@ 4.2 No Acyclicity Assumption
Conditional independencies arising from cyclic SCMs are entailed by the cyclic graphs [80]. It can be shown that, in general, there is no DAG encoding the conditional independencies which hold in such SCM [64]. Nonetheless, cyclic SCMs are widely used to model systems with feedback, and are applied in sociology, economics and biology, making this class of models a relevant target of interest for causal discovery techniques.
To test for such independencies, d-separation can be adapted to the cyclic setting under the assumption of causal sufficiency [101]. In causally insufficient scenarios, d-separation can be replaced with $\sigma$-separation [24, 25] applied to directed mixed graphs (DMGs), i.e. mixed graph (Subsection 2.24) without undirected edges.
@@
  • a collider $A*\!\!\rightarrow B\leftarrow\!\!*C$ where $B\not\in An(\textbf{Z})$, or
  • a non-collider $A\leftarrow B*\!\!-\!\!*C$ (or $A*\!\!-\!\!*B\rightarrow C$) where $B\in An(\textbf{Z})$ and $A$ (respectively $C$) is part of $SCC(B)$ (Equation 4.1).
@@
Cyclic Causal Discovery (CCD)
The Cyclic Causal Discovery (CCD) algorithm [80] has been the only provably sound (Subsection 2.12) approach to general directed graphs until the development of LiNG [48]. CCD is a constraint-based algorithm that follows the same initial procedure as the one of the PC algorthm, with five different orientation rules. CCD outputs a PAG $G$ which differ from the output of FCI for a couple of additional patterns:

### v2

@@ 2.7 Adding Prior Knowledge
A knowledge base $K$ is defined as an ordered pair $(\textbf{R},\textbf{F})$, where R is the set of required directed edges, while F is the set of forbidden directed edges.
@@ 3.1 Constraint-based Algorithms
The null $H_{0}$ and alternative hypotheses $H_{1}$ defined as $H_{0}:X\perp\!\!\!\perp_{P}Y\,|\,\textbf{Z}$ and $H_{1}:X\not\perp\!\!\!\perp_{P}Y\,|\,\textbf{Z}$, let $I(X,Y|\textbf{Z})$ denote a conditional independence (CI) test. The null hypothesis $H_{0}$ is not rejected if and only if the estimated p-value $\hat{I}(X,Y|\textbf{Z})$ is higher than a chosen significance level $\alpha$:
@@
where Z is a subset of V.
[…]
@@
Conditional Independence and Data Types
Constraint-based techniques are essentially agnostic of the specific conditional independence test that is being used. Indeed, it is possible to take advantage of such approaches in a wide variety of scenarios, as long as the assumptions of the said test are satisfied. While the main focus of causal discovery studies has been into either discrete or continuous settings, recent advances in conditional independence testing [61, 62] extend existing tests to mixed-data.
3.1.1 Peter-Clark (PC)
One of the most studied algorithm that leverages the CISs is the Peter-Clark (PC) algorithm [7] together with its variants [39, 63].
The first step of the procedure consists in defining a complete undirected graph over the variables of the given dataset D. Subsequently, a sequence of CI tests are performed following an heuristic strategy [39], in order to minimize the number of tests needed. For instance, it is known that the power of CI test decreases when the size of the conditioning set increases [64], due to the curse of dimensionality. A common approach consists in selecting an upper limit to the size of the conditioning set, discarding computational-intensive time-wasting tests with low significance levels.
The obtained independence statements are then used to remove the associated edges and to identify the underlying skeleton. Finally, the remaining edges are oriented according to a set of rules [31] that leverage the identified v-structures and acyclicity property.
@@ 3.1.1 Peter-Clark (PC)
3.1.2 Fast Causal Inference (FCI)
A first extension of the PC algorithm to the causal insufficient setting (Subsection 2.6) is represented by the Fast Causal Inference (FCI) algorithm [65, 34]. Specifically, the FCI algorithm relaxes both the assumption of no latent confounding [6] and no selection bias [66] in the observational setting, pushing the causal discovery problem a step closer to real-world scenarios. In this context, the authors leverage the definition of discriminating path to derive a new set of orientation rules.
@@ 3.1.2 Fast Causal Inference (FCI)
Let $G$ be an ancestral graph, a path $\pi=(X-\dots-W-Z-Y)$ between $X$ and $Y$ is a discriminating path for $Z$ if i) $\pi$ contains at least three edges, ii) $X$ is not adjacent to $Y$, iii) $Z$ is adjacent to $Y$, and iv) every vertex between $X$ and $Z$ is a collider on $\pi$ and parent of $Y$.
@@ 3.2 Score-based Algorithms
In the next few paragraphs, a set of properties for scoring criteria are introduced, before shifting the focus on an optimal two-step procedure for the causal sufficient scenario (Definition 2.6).
@@
Let D be a dataset associated with a probability distribution $P$, and let $G$ and $H$ be two graphs. A scoring criterion $\mathcal{S}$ is said to be consistent in the limit of the number of samples if and only if:
@@
  • $X\not\perp\!\!\!\perp_{P}Y\,|\,Pa(X)\implies\mathcal{S}(H,\textbf{D})>\mathcal{S}(G,\textbf{D})$, and
  • $X\perp\!\!\!\perp_{P}Y\,|\,Pa(X)\implies\mathcal{S}(H,\textbf{D})<\mathcal{S}(G,\textbf{D})$.
@@
This property guarantees that any deletion of an unnecessary edge will produce a higher score value, allowing the definition of an optimal greedy search algorithm.
One of the most commonly used score criterion is the Akaike Information Criterion (AIC) [68]:
$AIC=2k-2\ln{\hat{L}}$ (3.5)
where $k$ is the number of parameters of the model and $\hat{L}$ is the maximum value of the likelihood for the given model. Models achieving a lower value of AIC are preferred, i.e. they explain better the observed data. Another common scoring criterion is offered by the Bayesian Information Criterion (BIC) [69], also known as the Schwarz Information Criterion:
$BIC=k\ln{n}-2\ln{\hat{L}}$ (3.6)
which differs from AIC due to the parameters penalty term that takes into account the number of observations $n$. Others commonly used scoring critera are the Bayesian Dirichlet equivalent uniform (BDeu) [70] and the Bayesian Dirichlet sparse (BDs) [71].
@@
Let $G]^{*}$ be the equivalence class that is a perfect map of the probability distribution $P$ and D the associated dataset. $G]^{*}$ is said to be the optimal equivalence class if and only if:
@@
in the limit of the number of samples, for any consistent scoring criterion $\mathcal{S}$, following Definition 3.6.
3.2.1 Greedy Equivalent Search (GES)
The Greedy Equivalence Search (GES) [40, 72] is optimal in the limit of the number of samples [67]. The first step of the algorithm consists in the initialization of the empty graph $G$. The algorithm is composed by two phases: the forward search and the backward search. In the forward search phase, i) $G$ is modified by repeatedly adding the edge that has the highest difference in score (i.e. delta score), until there is no such edge that increases the score. In the backward search phase, ii) the edge that again achieves the highest delta score is repeatedly removed. The algorithm terminates once it reaches a local maximum during the backward search phase.
@@ 3.2.1 Greedy Equivalent Search (GES)
3.2.2 Fast GES (FGES)
Score-based algorithms are as fast as the computation of the chosen scoring criterion is. Leveraging the properties of the score function, it is possible to minimize the number of computations needed by storing previous intermediate evaluations. Not only this optimization reduces the computation time considerably, but also allows the application of these methods to high-dimensional datasets [61, 41]. This “fast” variant of GES (FGES) caches partial graph scores, significantly increasing the memory usage, since relevant fragments of the graph may be considered. Moreover, computationally expensive sections of the algorithm can be parallelized, taking advantage of high performance computing (HPC) settings.
@@ 3.3 Hybrid Algorithms
3.3.1 Adaptively Restricted GES (ARGES)
@@
The novelty of this hybrid version of GES stems from the concept of admissible edge. Let $G$ be a CPDAG and $X$ and $Y$ be a pair of non adjacent vertices on it. Adding an edge between $X$ and $Y$ is admissible for the graph $G$ if i) $X$ and $Y$ are adjacent in the (estimated) skeleton of $G$, or ii) there exists a node $Z$ such that $X\rightarrow Z\leftarrow Y$ is a v-structure in $G$.
@@
3.3.2 Greedy FCI (GFCI)
Score-based causal discovery algorithms such as GES and FGES are asymptotically correct, but are not designed to work in a causal insufficient scenario (Definition 2.6), where unmeasured confounders are present in the true graph. Constraint-based causal search algorithms, such as FCI, are asymptotically correct even with unmeasured confounders, but often perform poorly on small samples. The Greedy Fast Causal Inference (GFCI) [43] algorithm combines score-based and constraint-based algorithms improving over the previous results while being asymptotically correct under causal insufficiency.
@@ 3.4 Other Methods
3.4.1 Hidden Compact Representation (HCR)
@@
The authors address the problem of identifiability (Definition 2.14) of the model and prove that, under mild conditions, the causal graph recovered from observational data is identifiable. The method is tested against both synthetic and real-world data, providing reference values for performance metrics. In these experiments, HCR outperforms linear models in terms of accuracy and sensitivity, especially when the additive noise assumption does not hold.
3.4.2 Bivariate Quantile Causal Discovery (bQCD)
The bivariate quantile causal discovery (bQCD) [45] technique is designed to uncover cause-effect pairs in the bivariate setting. By re-expressing independence statements in light of the minimum description length (MDL) [73], the authors build a discovery procedure by using quantile scoring.
Following [74], let $X$ and $Y$ be two random variables with joint, marginal and conditional distributions denoted by $F_{XY}$, $F_{X}$ and $F_{X|Y}$, respectively. The key concept here is that a lower complexity follows from a correct causal orientation of the $(X,Y)$ pair, since it is a more informative representation of the associated data.
Hence, the Kolmogorov complexity $K(F)$ is defined as the length of the shortest program $F$ that outputs $F(X)$. Since $K(F)$ measures the information contained in $F$, authors in [75] state that if $X$ causes $Y$, then $K(F_{X})+K(F_{Y|X})\leq K(F_{Y})+K(F_{X|Y})$. The problem is that $K(F)$ cannot be computed in practice. Therefore, the authors rely on the MDL principle as a proxy for the Kolmogorov complexity. Such an approximation can be performed by estimating the population quantiles through nonparametric quantile regression.
The resulting procedure is robust to outliers and can be generalized to a wide range of distributions, although it requires that all population quantiles are computable, which could be a limiting factor in real-world applications.
3.4.3 Linear Non-Gaussian Acyclic Model (LiNGAM)
@@
where B is matrix consisting of the coefficients of the variables in the associated set of equations, x is the vector of variables and e is the vector of noises.
@@
LiNGAM [23, 38] relies on Independent Component Analysis (ICA) [76] to identify a possible solution for A. In fact, multiple mixing matrices A are feasible solutions for the given joint probability distribution. This technique is essentially focused on discovering asymmetries in the sample distribution to determine the correct causal ordering. Once such ordering has been discovered, the causal graph is built by recovering all and only the edges coherent with the order.
LiNGAM has been extended later for causally insufficient settings [46]. Let f be the vector of latent variables and $\boldsymbol{\Lambda}$ the matrix of the connections strength between f and x, then:
@@
LiNGAM consistently estimates the connection matrix B. While standard ICA does not scale well in high-dimensional settings, approximated variants of ICA can be used to compute the components with a predefined fix number of iterations with reasonable precision. This leads to an efficient solution in presence of non-Gaussian noise and causally insufficient datasets.
3.4.4 Continuous Optimization (NOTEARS)
In the “DAGs with NO TEARS” [47] algorithm, the causal discovery problem is reduced to a continuous optimization problem. The acyclicity constraint is expressed as an equality constraint $h(\textbf{W})=0$, where $h$ is a smooth differentiable function that measures the “DAG-ness” (i.e. a quantification of the acyclicity violations) of a given adjacency matrix W:
@@ 3.4.4 Continuous Optimization (NOTEARS)
Coefficients in W smaller than a fixed threshold $\omega>0$ are set to zero, rounding the solution with an arbitrary precision. The evaluation of the matrix exponential is $O(n^{3})$, i.e. cubic in the number of vertices. Given the low computational complexity, NOTEARS outperforms existing methods when both the in-degree and the sample size are large. However, even if the authors compared their approach to other methods (e.g. PC, FGES, LiNGAM), this algorithm is focused on exploring the DAG space, rather than addressing Definition 2.9 directly, i.e. an acyclic graph may not necessarily be a causal graph.
3.5 Comparison between Methods
Now that different classes of methods have been explored, one might be tempted to ask which approach performs better or is faster. Unfortunately, it is difficult to restrict the focus on a specific algorithm due to both theoretical and practical limitations, but still, one could at least try to compare advantages and limitations across classes of algorithms.
On the theoretical side, only a few contributions [7, 47] provide proofs of the asymptotic time complexity of such algorithms, making the comparison impossible due to the lack of references. On the practical side, there exist cases in which statistical criteria enforced during the evaluation steps of different algorithms do not match [77, 78], and others where computational time is recorded using absolute measurements [78, 79], i.e. seconds. These approaches not only make the comparison across different hardware configurations unfair, but are also affected by the size of the input. Furthermore, these issues are exacerbated by the fragmentation of the software packages across multiple programming languages, which are known to have different degrees of efficiency.
Nonetheless, a recent contribution [80] addresses these experimental pitfalls and conducts a rigorous analysis based on both synthetic and real-world data. While the authors do not exclude the existence of others potentially confounding factors, they essentially conclude that there is no systematic difference across classes of algorithms in terms of accuracy and speed, even taking into account a wide range of scenarios. Other works [81, 60] state that constraint-based algorithms are fast in general, but early mistakes undermine the construction of the final structure due to the chained evaluation of conditional independence tests, especially in high dimensional settings.
These last considerations seem to push towards score-based approaches, also considering the possibility of leveraging parallelization during the score evaluation, but further experiments are needed [80] in order to take into account others methods properly.
@@ 4.1 Cyclic SCM
In a SCM, the causal graph induces a functional set F where equations follow the decomposition enforced by the causal edge assumption, Subsection 2.9. If the causal graph is acyclic, then the SCM itself is called acyclic, or recursive SEM. The concept of recursion is linked to the hierarchical order that arises from the topological ordering of the underlying DAG. Indeed, it is possible to define a sequence $X_{1},X_{2},\dots,X_{n}$ of vertices over V such that for any $X_{i}$ and $X_{j}$ where $i<j$, $X_{j}$ is not a cause of $X_{i}$ [82].
@@ 4.2 No Acyclicity Assumption
Conditional independencies arising from cyclic SCMs are entailed by the cyclic graphs [48]. It can be shown that, in general, there is no DAG encoding the conditional independencies that hold in such SCM [83]. Nonetheless, cyclic SCMs are widely used to model systems with feedback, and are applied in sociology, economics and biology, making this class of models a relevant target of interest for causal discovery techniques.
To test for such independencies, d-separation can be adapted to the cyclic setting under the assumption of causal sufficiency (Definition 2.6) [84]. In causally insufficient scenarios, d-separation can be replaced with $\sigma$-separation [52, 33] applied to directed mixed graphs (DMGs), i.e. mixed graph (Subsection 2.25) without undirected edges.
@@
  • a collider $X*\!\!\rightarrow Y\leftarrow\!\!*Z$ where $Y\not\in An(\textbf{Z})$, or
  • a non-collider $X\leftarrow Y*\!\!-\!\!*Z$ (or $X*\!\!-\!\!*Y\rightarrow Z$) where $Y\in An(\textbf{Z})$ and $X$ (respectively $Z$) is part of $SCC(Y)$ (Equation 4.1).
@@
4.2.1 Cyclic Causal Discovery (CCD)
The Cyclic Causal Discovery (CCD) algorithm [48] has been the only provably sound (Subsection 2.12) approach to general directed graphs until the development of LiNG [49] (Subsection 4.2.2). CCD is a constraint-based algorithm that follows the same initial procedure as the one of the PC algorithm, with five different orientation rules. CCD outputs a PAG $G$ that differs from the output of FCI for a couple of additional patterns: